This is the FOURTH time I'm posting this question please post the full answer of ALL parts A,B,C,D,E. If you can't ,allow somebody 3. Two firms produce luxury sheepskin auto seat covers, Western Where (WW) and B.B.B. Sheep (BBBS). Each firm has a cost funcelse to do it.Thank you!
Answer #1

Page 30o-6 30 210x30)-(30)2-1-5(30)2 6300-100-1350 4050 Ao3 = 00-3)9301-1-542 16p5%)-35%-392-1642 2. Similarly BR2(%) : 1928B820 9 3037-5-억59:345-2848-125 Doveyeterom Max Ti St.92A -le -39-Aqo-3a1)|1-309-1。592 = 1500-39-90 Hilar 35% (-54 (180-29)9-1-532 180%-3.542 =Page ond fom BBB -4억:14 = 300-36 158-58- -330406-991-5-2013 An (158-6-50x alot 2-(105) (91。42)2 -3454.612-688-2346 = a066-58